# Disentangling representations of retinal images with generative models

### v2

@@ Abstract
Retinal fundus images play a crucial role in the early detection of eye diseases. However, the impact of technical factors on these images can pose challenges for reliable AI applications in ophthalmology. For example, large fundus cohorts are often confounded by factors like camera type, bearing the risk of learning shortcuts rather than the causal relationships behind the image generation process. Here, we introduce a population model for retinal fundus images that effectively disentangles patient attributes from camera effects, enabling controllable and highly realistic image generation. To achieve this, we propose a disentanglement loss based on distance correlation. Through qualitative and quantitative analyses, we show that our models encode desired information in disentangled subspaces and enable controllable image generation based on the learned subspaces, demonstrating the effectiveness of our disentanglement loss. The project’s code is publicly available [^1] https://github.com/berenslab/disentangling-retinal-images.
@@ 1 Introduction
Retinal fundus images are medical images that capture the back of the eye and show the retina, optic disc, macula, and blood vessels. Due to their nature as color photographs acquired through the pupil, they are inexpensive, non-invasive, and quick to obtain, making them available even in resource-constrained regions. Despite their high availability, fundus images can be used to detect the presence not only of eye diseases, but also cardiovascular risk factors or neurological disorders using deep learning (Poplin et al., 2018; Kim et al., 2020; Rim et al., 2020; Son et al., 2020; Sabanayagam et al., 2020; Xiao et al., 2021; Rim et al., 2021; Nusinovici et al., 2022; Cheung et al., 2022; Tseng et al., 2023).
@@
However, deep learning models require large datasets, and large medical imaging cohorts are often heterogeneous, with technical confounding being ubiquitous. For fundus images, variations in camera types, pupil dilation, image quality, and illumination levels can affect image generation. Additionally, biases due to different hospital sites and patient study selection can lead to spurious feature correlations of these factors with biological variations in the data. Deep learning models run the risk of learning shortcuts based on such spurious correlations (Geirhos et al., 2020), instead of specific patient-related image features (Fig. 1a). As a result, they may only perform well within the same distribution as the training data and latent space analyses may be misleading (Müller et al., 2022). For example, a retinal imaging dataset could be collected from two hospitals using cameras from different manufacturers (Fay et al., 2023). Camera A produces images appearing different in hue from camera B. The first hospital may treat predominantly Latin American patients, while the second hospital may treat predominantly Caucasians. In this scenario, it is simpler for a deep learning model to infer a patient’s ethnicity from a fundus image by recognizing hue rather than understanding the hidden causal relationship between ethnicity and phenotypic image features (Castro et al., 2020).
@@
One solution for dealing with confounders in data is subspace learning, which combines representation learning and disentanglement. Latent subspaces can be constructed from a causal graph using domain knowledge about what influences the image generation process (Castro et al., 2020). In our case, we assume a simplified causal model that splits the fundus image generation into patient characteristics and technical factors (Fig. 1a). We assume a shortcut connection due to spurious correlations between a patient attribute, such as ethnicity, and the camera as a technical factor (Fig. 1a, blue dashed arrow). To prevent shortcut learning, we aim to disentangle these factors of variation by encoding them into statistically independent latent subspaces (Bengio et al., 2013; Higgins et al., 2018).
@@
One common strategy for acquiring disentangled representations makes use of generative image models such as variational autoencoders (VAEs) (Kingma and Welling, 2014). Generative models offer a valuable inductive bias for representation learning by reconstructing images from the latent space and can be extended to learn independent subspaces (Fig. 1b) (Higgins et al., 2017; Klys et al., 2018; Tschannen et al., 2018). Previous methods primarily focused on achieving a disentangled latent space, often neglecting the simultaneous pursuit of high-quality image generation. However, when latent space disentanglement and image generation are addressed simultaneously, generative image models provide a built-in interpretation method through controllable feature changes in the image space (Fig. 1d).
@@
Figure 1: Disentangling representations of retinal images with generative models. In panel a, we present our conceptual causal graph outlining the image generation process (Fay et al., 2023). Within this framework, we break down retinal fundus image generation into patient-specific and technical factors of variation and highlight a potential shortcut connection for deep learning models. In panel b, we introduce a strategy to prevent shortcut learning by using a generative model as an independent subspace learner. Panel c shows our metric for disentanglement: an ideal confusion matrix between the available labels (columns) and the learned subspaces (rows). Panel d shows controllable fundus image generation from swapped ethnicity subspaces.
@@ 2 Related work
In medical imaging, disentangled representations are relevant to many research questions related to image synthesis, segmentation, registration, and causal- or federated learning, e.g., for disease decomposition, harmonization, or controllable synthesis Liu et al. (2022). For example, Fay et al. (2023) present a predictive approach to avoid shortcut learning for MRI data from different scanners. To minimize the dependencies between a task-specific factor (e.g. patient attribute) and a confounding correlated factor (e.g. scanner), they estimate a lower bound for MI. However, as mentioned above, a lower bound MI estimate is inconsistent with MI minimization tasks. In ICAM (Bass et al., 2020, 2021), the authors build on a VAE-GAN approach for image-to-image translation (Lee et al., 2018) to disentangle disease-relevant from disease-irrelevant features. However, disentanglement in ICAM requires longitudinal data, where patient images are available at different disease stages. In a similar approach, Ouyang et al. (2021) embed MRI images into independent anatomical and modality subspaces by image-to-image translation. Like ICAM, they have the data constraint that their method only works with images of the same patient from different modalities. In another attempt to explain classifier features, StylEx (Lang et al., 2021) integrates a classifier into StyleGAN training. They have a classifier-specific intermediate latent space and can visualize image features that are important for the classifier decision. Although the authors apply their approach to the retinal fundus domain, they do not address spurious correlation and shortcut learning.
@@ 3 Methods
This section introduces the proposed extensions to the generative model and training constraints to disentangle the representations, prevent shortcut learning, and still enable controllable high-resolution image generation. First, we introduce the dataset, an encoder model and our disentanglement loss as preliminary steps towards the complete image generation pipeline.
@@ 3.1 Data
We trained our models on retinal fundus images provided by EyePACS Inc. (Cuadros and Bresnick, 2009; Eye, 2008), an adaptable telemedicine system for diabetic retinopathy screening from California. All data was anonymized by the data provider. We filtered for healthy fundus images with no reported eye disease which were labelled as “good” or “excellent” quality. This resulted in 75,989 macula-centered retinal fundus images of 24,336 individual patients. For training, we split the data by patient-identity into 60% training, 20% validation, and 20% test sets.
@@ 3.2 Encoder model for mapping factors of variation
$(\theta^{*},\psi^{*})$ $=\operatorname*{arg\,min}_{\theta,\psi}\,L_{\text{CE}}(\theta,\psi)$ (1)
@@ 3.3 Disentanglement loss
$(\theta^{*},\psi^{*})$ $=\operatorname*{arg\,min}_{\theta,\psi}\,L_{\text{CE}}(\theta,\psi)+\lambda_{\text{DC}}L_{\text{DC}}(\theta)$ (3)
@@
We use the empirical distance correlation (Székely et al., 2007) to measure correlation between batch samples of subspace tensors:
@@ 3.4 Generative image model
For the generative model, the overall goal was to generate realistic fundus images from a disentangled latent space. We considered the image generation task by learning the mapping between two domains $M:\mathcal{X}\rightarrow\mathcal{\hat{X}}$ with a latent space model. We split $M$ into two components, $M=d\circ g$, where the encoder/discriminator $d$ maps to an intermediate latent representation $d:\mathcal{X}\rightarrow\mathcal{W}$ and the generator $g:\mathcal{W}\rightarrow\mathcal{\hat{X}}$ maps to the output image domain.
We chose to work with the StyleGAN2 architecture (Karras et al., 2020), a state-of-the-art generative adversarial network for high-fidelity and high-resolution images. The StyleGAN architecture has two features that makes it a potentially good candidate for disentanglement: (1) it consists of a mapping network $m:\mathcal{Z}\rightarrow\mathcal{W}$ that maps a latent vector into an intermediate space, which then controls the styles in each convolutional layer in the generator with adaptive instance normalization, and (2) it can control scale-specific image generation by a progressively growing generator.
@@
As GANs were not originally designed for representation learning, they do not provide an encoder model. Therefore, we extended the StyleGAN2 architecture to embed real images in the latest space for representation learning. Inspired by the work of invGAN (Ghosh et al., 2022) for GAN-inversion, we also trained the discriminator as an encoder. We added two independent, fully connected layers and returned an estimate of the latent representation $\hat{w}$ for our first inversion loss
@@
The main methodological contribution of this work is to extend the StyleGAN2 architecture (with GAN-inversion for representation learning) to an independent subspace learner. Therefore, we modified the architecture to sample individual latent codes from a standard normal distribution $z_{k}\sim\mathcal{N}(0,I_{d_{k}})$ and learned individual 8-layer mapping networks for each subspace $m_{\eta}^{k}:\mathcal{Z}_{k}\rightarrow\mathcal{W}_{k}$ (Fig. 4, left), where each subspace can have a different dimensionality $w_{k}\in\mathbb{R}^{d_{k}}$. In setting up the generative image model, we assumed that we do not know all the underlying factors of variation or have access to their labels. Therefore, we provided the model with a free patient-identity subspace $w_{3}$ (Fig. 4), in which the encoder/discriminator can capture further information required to reconstruct high-resolution retinal fundus images. As in the encoder model (Sec. 3.2), we encoded label information for real images by training linear classification heads $C_{\psi_{k}}$ for the first two subspaces (Fig. 4, right). To avoid subspace correlation, we penalized the classification task with the mean of all distance correlation measures between subspaces $\hat{w}_{k}$ that the discriminator outputs.
@@
$(\tau^{*},\psi^{*})$ $=\operatorname*{arg\,min}_{\tau,\psi}\,-L_{\text{GAN}}(\eta^{*},\tau)+\lambda_{w}L_{w}(\eta^{*},\tau)+\lambda_{p}L_{p}(\eta^{*},\tau)+\lambda_{\text{C}}L_{\text{CE}}(\tau,\psi)+\lambda_{\text{DC}}L_{\text{DC}}(\tau)$ (14)
@@
In practice, we optimized the objectives for the generator (Eq. 13) and the discriminator (Eq. 14) with batch estimates and stochastic gradient descent. Moreover, we used separate optimizers for generator and discriminator optimization. Appendix C contains additional information about the generative model training.
Figure 5: Encoder disentanglement performance for age-camera setup. We trained encoder models on retinal images, where the first subspace $w_{\text{age}}$ encoded age $y_{\text{age}}$, and the second subspace $w_{\text{cam}}$ encoded camera information $y_{\text{cam}}$. In panels a and b, we show the accuracy improvement over chance level accuracy. As a baseline model in a, we only trained linear classifiers on top of subspaces. In b, we further disentangle the subspaces using our distance correlation loss. Panel c shows the distance correlation measure between subspaces for both the baseline method and our proposed method for subspace disentanglement. In panels d and e, we provide t-SNE visualizations comparing a baseline model with a model with disentanglement loss, respectively.
@@ 4 Results
To evaluate the success, we need to define a quantitative measure of disentanglement. While there is no universally accepted definition of disentanglement, most agree on two main points (Eastwood and Williams, 2018; Carbonneau et al., 2022): (1) factor independence and (2) information content. Factor independence means that a factor affects only a subset of the representation space, and only this factor affects this subspace. We refer to this property as modularity. As another criterion for factor independence some argue that the affected representation space by a factor should be as small as possible, a principle known as compactness. The second agreed-upon criterion is the information content or explicitness. In essence, a disentangled representation should completely describe all factors of interest. To assess disentanglement, three distinct families of metrics have been proposed: intervention-based, predictor-based, and information-based metrics (Carbonneau et al., 2022).
In our case, we opted for a predictor-based metric to evaluate both modularity and explicitness, excluding compactness because we fixed the subspace dimensions. For our predictor, we employed a k-nearest neighbors (kNN) classifier with k=30. Using this classifier, we generated a confusion matrix, comparing subspaces with associated labels, for example ethnicity and camera (Fig. 6). Given the presence of unbalanced classes, we report the accuracy improvement over chance level accuracy. For the confusion matrix, modularity is evident by a high classification accuracy improvement for one factor (e.g. ethnicity) in only one subspace. Explicitness, on the other hand, is evident when the accuracy improvement is high across the entire matrix. Consequently, when both modularity and explicitness are effectively addressed, the confusion matrix has high values along the main diagonal as well as low values on the off-diagonals (Fig. 6).
Figure 6: Example of a confusion matrix to measure disentanglement in terms of modularity and explicitness.
For both the predictive and the generative task, we conducted two experiments, configuring the model to learn specific subspaces. In the initial experiment, the first subspace $w_{\text{age}}\in\mathbb{R}^{4}$ encoded age information, while the second subspace $w_{\text{cam}}\in\mathbb{R}^{12}$ captured camera-related details. In the second experiment, we designed the first subspace $w_{\text{ethn}}\in\mathbb{R}^{8}$ to encode ethnicity information, while the second subspace focused again on camera-related features $w_{\text{cam}}\in\mathbb{R}^{12}$. The dimensionality of each subspace was selected based on the dimensions of the labels we aimed to encode.

### v3

@@ Abstract
Retinal fundus images play a crucial role in the early detection of eye diseases. However, the impact of technical factors on these images can pose challenges for reliable AI applications in ophthalmology. For example, large fundus cohorts are often confounded by factors such as camera type, bearing the risk of learning shortcuts rather than the causal relationships behind the image generation process. Here, we introduce a population model for retinal fundus images that effectively disentangles patient attributes from camera effects, enabling controllable and highly realistic image generation. To achieve this, we propose a disentanglement loss based on distance correlation. Through qualitative and quantitative analyses, we show that our models encode desired information in disentangled subspaces and enable controllable image generation based on the learned subspaces, demonstrating the effectiveness of our disentanglement loss. The project code is publicly available [^1] https://github.com/berenslab/disentangling-retinal-images.
@@ 1 Introduction
Figure 1: Disentangling representations of retinal images with generative models. In panel a, we present our conceptual causal graph outlining the image generation process. Within this framework, we break down retinal fundus image generation into the causal variables patient attribute (B), camera (C), and style (D) and highlight an unknown confounder (A), which can lead to shortcut connections for deep learning models. In panel b, we introduce a strategy to prevent shortcut learning by using a generative model as an independent subspace learner. Panel c shows our metric for disentanglement: an ideal confusion matrix between the available labels (columns) and the learned subspaces (rows). Panel d shows controllable fundus image generation from swapped ethnicity subspaces.
@@
Retinal fundus images are medical images that capture the back of the eye and show the retina, optic disc, macula, and blood vessels. Due to their nature as color photographs acquired through the pupil, they are inexpensive, non-invasive, and quick to obtain, making them available even in resource-constrained regions. Despite their high availability, fundus images can be used to detect the presence of not only eye diseases, but also cardiovascular risk factors or neurological disorders using deep learning (Poplin et al., 2018; Kim et al., 2020; Rim et al., 2020; Son et al., 2020; Sabanayagam et al., 2020; Xiao et al., 2021; Rim et al., 2021; Nusinovici et al., 2022; Cheung et al., 2022; Tseng et al., 2023).
@@
However, deep learning models require large datasets, and large medical imaging cohorts are often heterogeneous, with technical confounding being ubiquitous. For fundus images, variations in camera types, pupil dilation, image quality, and illumination levels can affect image generation. Additionally, biases due to different hospital sites and patient study selection can lead to spurious feature correlations of these factors with biological variations in the data. Deep learning models run the risk of learning shortcuts based on such spurious correlations (Geirhos et al., 2020), instead of specific patient-related image features (Fig. 1a). As a result, they may only perform well within the same distribution as training data and latent space analyses can be misleading (Müller et al., 2022). For example, a retinal imaging dataset could be collected from two hospitals using cameras from different manufacturers (Fay et al., 2023). Camera A produces images that appear different in hue from camera B. The first hospital may treat predominantly Latin American patients, while the second hospital may treat predominantly Caucasians. In this scenario, it is simpler for a deep learning model to infer a patient’s ethnicity from a fundus image by recognizing the hue rather than understanding the hidden causal relationship between ethnicity and phenotypic image features (Castro et al., 2020).
@@
One approach to address data confounders is subspace learning, which combines representation learning with disentanglement. In our work, we constructed latent subspaces based on a causal graph, leveraging domain knowledge about the factors influencing the image generation process (Castro et al., 2020). We used a simplified causal model that divides fundus image generation into three causal variables: patient attribute (B), camera (C), and style (D) (Fig. 1a). Style features in the literature of causal representation learning and image generation are attributed to information which is agnostic to content features (B and C in our case) (Yao et al., 2024b; Havaei et al., 2021). Due to the presence of an unknown confounder (A), we assume that there is a statistical association between B, C, and D, even though they are not directly causally related. This association could lead to shortcut connections driven by spurious correlations, particularly between factors like patient attributes (e.g., ethnicity) and technical factors like the camera (Fig. 1a). To prevent shortcut learning, we disentangle these factors of variation by encoding them into statistically independent latent subspaces (Bengio et al., 2013; Higgins et al., 2018).
@@
One common strategy for acquiring disentangled representations is to use generative image models such as variational autoencoders (VAEs) (Kingma and Welling, 2014). Generative models offer a valuable inductive bias for representation learning by reconstructing images from the latent space and can be extended to learn independent subspaces (Fig. 1b) (Higgins et al., 2017; Klys et al., 2018; Tschannen et al., 2018). Previous methods primarily focused on achieving a disentangled latent space, often neglecting the simultaneous pursuit of high-quality image generation. However, when latent space disentanglement and image generation are addressed simultaneously, generative image models provide a built-in interpretation method through controllable feature changes in the image space (Fig. 1d).
@@
2 Causal background
Deep generative models have demonstrated great success in data generation and density estimation from observational data, but they face limitations such as lack of explainability, risk of learning spurious correlations and poor performance on out-of distribution data. To address these challenges, causal models offer several valuable benefits, such as distribution shift robustness, fairness, and interpretability (Schölkopf et al., 2021). The interpretability of causal models can be viewed in terms of both learning semantically meaningful representations of a system and controllably generating realistic counterfactual data. Causal models, particularly Structural Causal Models (SCMs), describe the data-generating processes and the complex causal relationships between variables, making them a natural complement to generative models (Schölkopf et al., 2021). Causal representation learning (CRL) focuses on identifying meaningful latent variables and their causal structures from high-dimensional data, ensuring the correct identification of the true generative factors (Schölkopf et al., 2021).
In our causal diagram the fundus image (X) is influenced by several factors: patient attribute (B), camera (C), and style (D). However, there is also an unknown confounder (A) that affects all of these factors and creates a statistical association between the causal variables B, C, D, even though they are not causally related (Castro et al., 2020; Schölkopf et al., 2021). We represent the causal relationships between these variables by a directed acyclic graph (DAG) which implies the factorization of causal conditionals (Schölkopf et al., 2021):
$P(X)=P(B|A)\cdot P(C|A)\cdot P(D|A).$
In theory, the confounder could be addressed by causal interventions, which is not feasible in our setting with only observational data. Therefore, in this work, we address the confounder by disentangling the causal variables B, C, D in the latent space of a generative model. Our approach is inspired by our DAG structure, which implies a disentangled factorization of causal conditionals due to the causal Markov condition (Castro et al., 2020). Regarding CRL, our causal variables B and C are identifiable because of the invariance principle introduced in (Yao et al., 2024a). For these variables, we use labels together with linear classifiers to enforce content encoding in latent subspaces. Intuitively, with our subspace classifiers we force images with the same label to have the same content latent subspaces. Identifying the style variable (D) is more complicated, but we can use Corollary 3.9-3.11 from (Yao et al., 2024b) to identify style, assuming that it is independent of the content variables B and C.
@@ 3 Related work
In medical imaging, disentangled representations are relevant to many research questions related to image synthesis, segmentation, registration, and causal- or federated learning, e.g., for disease decomposition, harmonization, or controllable synthesis Liu et al. (2022). For example, Fay et al. (2023) present a predictive approach to avoid shortcut learning for MRI data from different scanners. To minimize the dependencies between a task-specific factor (e.g. patient attribute) and a confounding correlated factor (e.g. scanner), they estimate a lower bound for MI. However, as mentioned above, a lower bound MI estimate is inconsistent with MI minimization tasks. In ICAM (Bass et al., 2020, 2021), the authors build on a VAE-GAN approach for image-to-image translation (Lee et al., 2018) to disentangle disease-relevant from disease-irrelevant features. However, disentanglement in ICAM requires longitudinal data, where patient images are available at different disease stages. In a similar approach, Ouyang et al. (2021) embed MRI images into independent anatomical and modality subspaces by image-to-image translation. Like ICAM, they have the data constraint that their method only works with images of the same patient from different modalities. In another attempt to explain classifier features, StylEx (Lang et al., 2021) integrates a classifier into StyleGAN training. They have a classifier-specific intermediate latent space and can visualize image features that are important for the classifier decision. Although the authors apply their approach to the retinal fundus domain, they do not address potential spurious correlations with latent space disentanglement. In Galati et al. (2023), another work based on StyleGAN, they present a semi-supervised adaptation framework for brain vessel segmentation between two imaging domains: magnetic resonance angiography to venography. However, here they assume that all vessel properties are disentangled within the StyleGAN latent space and do not add any new regularizers regarding latent space disentanglement. In Havaei et al. (2021) they train a conditional adversarial generative model with two latent spaces that encode style and content which they also disentangle with two regularization steps. However, compared to our method, they only generate low resolution images and do not focus on the generation of high frequency details. In addition, their optimization procedure is very involved compared to ours, with two generators and five discriminators.
@@ 4 Methods
This section introduces the proposed model extensions and training constraints to disentangle representations, prevent shortcut learning, and still enable controllable high-resolution image generation. First, we introduce the dataset, an encoder model and our disentanglement loss as preliminary steps towards the complete image generation pipeline.
@@ 4.1 Data
We trained our models on retinal fundus images provided by EyePACS Inc. (Cuadros and Bresnick, 2009; Eye, 2008), an adaptable telemedicine system for diabetic retinopathy screening from California. All data was anonymized by the data provider. We filtered for healthy fundus images with no reported eye disease which were labelled as “good” or “excellent” quality. This resulted in 75,989 macula-centered retinal fundus images of 24,336 individual patients. For our experiments, we split the data by patient-identity into 60% training, 20% validation, and 20% test sets.
@@ 4.2 Encoder model for mapping factors of variation
$(\theta^{*},\psi^{*})$ $=\operatorname*{arg\,min}_{\theta,\psi}\,L_{\text{CE}}(\theta,\psi),$ (1)
@@ 4.3 Disentanglement loss
$(\theta^{*},\psi^{*})$ $=\operatorname*{arg\,min}_{\theta,\psi}\,L_{\text{CE}}(\theta,\psi)+\lambda_{\text{DC}}L_{\text{DC}}(\theta),$ (3)
@@
We used the empirical distance correlation (Székely et al., 2007) to measure correlation between batch samples of subspace tensors:
@@ 4.4 Generative image model
For the generative model, the overall goal was to generate realistic fundus images from a disentangled latent space. We considered the image generation task by learning the mapping between two domains $M:\mathcal{X}\rightarrow\mathcal{\hat{X}}$ with a latent space model. We split $M$ into two components, $M=d\circ g$, where the discriminator/encoder $d:\mathcal{X}\rightarrow\mathcal{W}$ maps from the image domain to a latent representation, and the generator $g:\mathcal{W}\rightarrow\mathcal{\hat{X}}$ maps from the latent representation to the image reconstruction.
As the backbone generative image model we selected the StyleGAN2 architecture (Karras et al., 2020), a state-of-the-art generative adversarial network for high-fidelity and high-resolution images. The StyleGAN architecture has two features that makes it a good candidate for disentanglement: (1) a mapping network $m:\mathcal{Z}\rightarrow\mathcal{W}$ that maps a latent vector into an intermediate space, which then controls the styles in each convolutional layer in the generator with adaptive instance normalization, and (2) scale-specific image generation by a progressively growing generator.
@@
As GANs were not originally designed for representation learning, they do not provide an encoder model. Therefore, we extended the StyleGAN2 architecture to embed real images in the latent space for representation learning. Inspired by the work of invGAN (Ghosh et al., 2022) for GAN-inversion, we also trained the discriminator as an encoder. We added a fully connected layer to estimate the latent representation $\hat{w}$ for our first inversion loss
@@
The main methodological contribution of this work is to extend the StyleGAN2 architecture (with GAN-inversion for representation learning) to an independent subspace learner. Therefore, we modified the architecture to sample individual latent codes from a standard normal distribution $z_{k}\sim\mathcal{N}(0,I_{d_{k}})$ and learned individual 8-layer mapping networks for each subspace $m_{\eta}^{k}:\mathcal{Z}_{k}\rightarrow\mathcal{W}_{k}$ (Fig. 4, left), where each subspace can have a different dimensionality $w_{k}\in\mathbb{R}^{d_{k}}$. In setting up the generative image model, we assumed that we do not know all the underlying factors of variation or have access to their labels. Therefore, we provided the model with a content agnostic style space $w_{3}$ (Fig. 4), in which the encoder/discriminator can capture further information required to reconstruct high-resolution retinal fundus images which is independent from the other content subspaces.
As in the encoder model (Sec. 4.2), we encoded label information for real images by training linear classification heads $C_{\psi_{k}}$ for the first two subspaces (Fig. 4, right). To avoid subspace correlation, we penalized the classification task with the mean of all distance correlation measures between subspaces $\hat{w}_{k}$ that the discriminator outputs.
@@
$(\tau^{*},\psi^{*})=$ $\operatorname*{arg\,min}_{\tau,\psi}\,-L_{\text{GAN}}(\eta^{*},\tau)+\lambda_{w}L_{w}(\eta^{*},\tau)+\lambda_{p}L_{p}(\eta^{*},\tau)+\frac{\lambda_{\text{C}}}{K-1}\sum_{k=1}^{K-1}L_{\text{CE}}(\tau,\psi_{k})+\lambda_{\text{DC}}L_{\text{DC}}(\tau)$ (14)
@@
In practice, we optimized the objectives for the generator (Eq. 13) and the discriminator (Eq. 14) with batch estimates and stochastic gradient descent. Moreover, we used separate optimizers for generator and discriminator optimization. Appendix C offers additional information about the generative model training.
@@ 5 Results
To evaluate the success, we defined a quantitative measure of disentanglement. While there is no universally accepted definition of disentanglement, most agree on two main points (Eastwood and Williams, 2018; Carbonneau et al., 2022): (1) factor independence and (2) information content. Factor independence means that a factor affects only a subset of the representation space, and only this factor affects this subspace. We refer to this property as modularity. As another criterion for factor independence some argue that the affected representation space by a factor should be as small as possible, a principle known as compactness. The second agreed-upon criterion is the information content or explicitness. In essence, a disentangled representation should completely describe all factors of interest. To assess disentanglement, three distinct families of metrics have been proposed: intervention-based, predictor-based, and information-based metrics (Carbonneau et al., 2022).
In our case, we opted for a predictor-based metric to evaluate both modularity and explicitness, excluding compactness because we fixed the subspace dimensions. For our predictor, we employed a k-nearest neighbors (kNN) classifier with k=30. Using this classifier, we generated a confusion matrix, comparing subspaces with associated labels, for example ethnicity and camera (Fig. 1c). Given the presence of unbalanced classes, we report the accuracy improvement over chance level accuracy. For the confusion matrix, modularity is evident by a high classification accuracy improvement for one factor (e.g. ethnicity) in only one subspace. Explicitness, on the other hand, is evident when the accuracy improvement is high across the entire matrix. Consequently, when both modularity and explicitness are effectively addressed, the confusion matrix has high values along the main diagonal as well as low values on the off-diagonals (Fig. 1c).
We conducted two experiments, configuring the model to learn specific subspaces. In the initial experiment, the first subspace $w_{\text{age}}\in\mathbb{R}^{4}$ encoded age information, while the second subspace $w_{\text{cam}}\in\mathbb{R}^{12}$ captured camera-related details. In the second experiment, we designed the first subspace $w_{\text{ethn}}\in\mathbb{R}^{8}$ to encode ethnicity information, while the second subspace focused again on camera-related features $w_{\text{cam}}\in\mathbb{R}^{12}$. The dimensionality of each subspace was selected based on the dimensions of the labels we aimed to encode.